OpenStudy (anonymous):

The density (mass divided by volume) of water is 1.00 g/cm^3. What is this value in kilograms per cubic meter?

OpenStudy (anonymous):

My work: (1.00 g/cm^3)(1 kg/1000 g)[(100 cm^3 / m^3] = 0.1 kg/m^3 But the answer is 1.00kg * 10^3/m^3. What am I doing wrong?

OpenStudy (ash2326):

1 m = 100 cm 1 m^3= (100 cm)^3 so 1m^3=(100)^3 cm^3

OpenStudy (ash2326):

Do you get this? use this in your work and you'll get the correct answer

OpenStudy (anonymous):

Thanks.